OpenStudy (anonymous):

how many solutions are there to the equation, 11x-99=11(x-9)??

OpenStudy (anonymous):

well, they are equal. 11(x-9)=11*x - 11*9=11x-99 therefore, 11x-99=11x-99 add 99 to both sides gives 11x=11x divide by 11 on both sides gives x=x. the answer is technically an infinite amount of solutions

OpenStudy (anonymous):

maybe there was something you didnt specify?